# Homeomorphism groups of basilica, rabbit and airplane Julia sets

### v1

@@ B.1 The rabbit Julia sets
Let $U_{1}$ be the Fatou component of $c_{1}$ and $U_{0}$ that of the critical point 00. For a component $U$ of $\mathring{K_{c}}$, its image $f(U)$ is also some component of $\mathring{K_{c}}$ (this follows from the fact that $f_{c}$ is an open map) and, if $U$ does not contain 00 (i.e., $U\neq U_{0}$), then $f_{c}$ restricts to a homeomorphism from $U$ to $f_{c}(U)$.
@@
The projection to any component is a cut point. Thus, if $x\in\partial U$ and $y\in\partial V$, where $U$ and $V$ are distinct components of $\mathring{K_{c_{n}}}$, then the projection of $x$ to $V$ is a cut point that separates $x$ from $y$ [DH84a, Exposé I, Proposition 4]. This proves point 2 of A
For $x\in K_{c_{n}}$, we denote by $\ell(x)$ the number of components $V$ of $\mathring{K_{c_{n}}}$ such that $V\cap[x,\alpha_{1}]\neq\emptyset$, where $[x,\alpha_{1}]$ is the legal arc between $x$ and $\alpha_{1}$ (legal arcs were introduced in [DH84a, Exposé II, §6] as arcs réglementaires). By construction of legal arcs, this number is constant on each component $U$ of $\mathring{K_{c}}$, so we denote by $\ell(U)$ this common number (which, a priori, could be infinite). We claim that, for any component $U$, $|\ell(f_{c}(U))-\ell(U)|\leq 1$. If the legal arc between $x\in U$ and $\alpha_{1}$ does not contain 00 then there is equality, since $f_{c}$ maps components to components and, by [DH84a, Exposé IV, Lemme 1], it maps legal arcs to legal arcs injectively if the legal arc does not contain the critical point in its interior. In case $[x,\alpha_{1}]$ contains 00, we have that $[x,\alpha_{1}]=[x,0]\cup[0,\alpha_{1}]$. If $[f_{c}(x),f_{c}(0)]=[f_{c}(x),c]$ contains $\alpha_{1}$, then $\ell(f_{c}(U))=\ell(U)-1$ because $[f_{c}(x),c]=[f_{c}(x),\alpha_{1}]\cup[\alpha_{1},c]$ and $c$ belongs to the component $U_{1}$ which contains $\alpha_{1}$ in its boundary. If instead $[f_{c}(x),c]$ does not contain $\alpha_{1}$, then $\ell(f_{c}(U))=\ell(U)$.
We know that $U_{0}$ contains a point of the attractive cycle, so for any component $U$, there is $n$ such that $f_{c}^{n}(U)=U_{0}$. In particular, $\ell(U)\leq n+1$. So, if $x$ lies on some circle of $J_{c}$, i.e., $x\in\partial U$ for some component $U$ of $\mathring{K_{c}}$, then $\ell(x)=\ell(U)$ or $\ell(x)=\ell(U)-1$, depending on whether $x$ is the projection of $\alpha_{1}$ on $U$ or not. If $x$ and $y$ belong to circles, then the collection of circles that separate $x$ from $y$ is included in the collection of circles that separate $x$ from $\alpha_{1}$ and $y$ from $\alpha_{1}$. This is easily seen using the fact that the legal hull of $x$, $y$ and $\alpha_{1}$ (introduced in [DH84a, Exposé II, §7] as enveloppe réglementaire) is a topological tree. This collection is finite, since $\ell(x),\ell(y)<\infty$. This proves point 3 of A.
@@ B.2 The airplane Julia set
Let $\alpha_{1}$ be the point in $J_{c}$ with the external rays $3/7$ and $4/7$. This is the projection of 00 on the component $U_{1}$ of $\mathring{K_{c}}$ that contains $c$. The point $\alpha_{1}$ belongs to the circle $\partial U_{1}$ and is a cut point, since it separates 00 and $c$. The set of its preimages is dense, which proves point 1 of B.
@@
The point $\alpha_{1}$ is mapped to the point with external arguments ±$\pm 1/7$ which is mapped to the point with external arguments ±$\pm 2/7$. This last point is mapped to $\alpha_{1}$, which is thus 3-periodic. Let $U_{0}$ be the component that contains 00 and let $U_{2}$ be the image of $U_{1}$. These components form a 3-cycle and $U_{0}$ lies between $U_{1}$ and $U_{2}$. For two components $U,V$ of $\mathring{K_{c}}$, we denote by $\delta(U,V)$ the number of components that the legal arc from any point in $U$ to any point in $V$ crosses minus 1 (this number can be infinite). Since $U_{0}$ is between $U_{1}$ and $U_{2}$, we have $\delta(U_{1},U_{2})=\delta(U_{1},U_{0})+\delta(U_{0},U_{2})$. Because of [DH84a, Exposé IV, Lemme 1], we have $\delta(U_{2},U_{1})=\delta(f_{c}(U_{1}),f_{c}(U_{0}))=\delta(U_{1},U_{0})$ and $\delta(U_{0},U_{1})=\delta(f_{c}(U_{2}),f_{c}(U_{0}))=\delta(U_{2},U_{0})$. Since all these numbers are at least $1$, they must all be infinite.
@@
Assume that two circles $C_{1},C_{2}$ have a common point. Let $V_{1},V_{2}$ be the two components of $\mathring{K_{c}}$ such that $C_{i}=\partial V_{i}$. There is a large enough $k\in\mathbf{N}$ such that $f_{c}^{k}$ maps $V_{i}$ to $U_{j_{i}}$ (for $i=1,2$ and $j_{i}\in\{0,1,2\}$) . By [DH84a, Exposé IV, Lemme 1], $f_{c}^{k}$ maps the legal arc between $V_{1}$ and $V_{2}$ to the legal arc between $U_{j_{1}}$ and $U_{j_{2}}$, so it maps the point shared by $V_{1}$ and $V_{2}$ to a point shared by two of the $U_{i}$’s. We have a contradiction, since the $U_{i}$’s have at least one circle that separates them (in fact, infinitely many). This proves point 3 of B.
@@
Assume that two circles $C_{1},C_{2}$ are not separated by a third one. Let $V_{i}$’s be components as above. The legal arc $A$ joining the centers does not contain 00 and the same is true for all $f_{c}^{n}(A)$. Let $n$ be a large enough number so that $f^{n}_{c}(V_{i})$’s are some $U_{j}$’s. The images of $V_{1}$ and $V_{2}$ are distinct and thus separated by infinitely many circles. This implies that $A$ meets infinitely many circles. A contradiction, which proves point 2 of B.
@@
It remains to prove that every cut point has order $2$. Since every cut point has order at least $2$, it suffices to prove there is no cut point of order at least $3$. By contradiction, let $x$ be such a point. Since $K_{c}$ is obtained by filling circles from $J_{c}$, the complement $K_{c}\setminus\{x\}$ has at least $3$ connected components. Choose components $V_{1},V_{2},V_{3}$ of $\mathring{K_{c}}$ lying in distinct components of $K_{c}\setminus\{x\}$. Let $T$ the legal hull of the centers of the $V_{i}$’s. Up to choosing a smaller tree, we may assume that $T$ does not contain 00 in the interior of any of its arcs. It is a tripod with $3$ arcs and $x$ as center. We define a sequence of legal trees in the following way. Let $T^{(0)}=T$ and assume that $T^{(n)}$ has been constructed and does not contain 00 in the interior of any of its arcs. Then, by [DH84a, Exposé IV, Lemme 1], $f_{c}$ maps $T^{(n)}$ injectively to $f_{c}(T^{(n)})$. If the latter does not contain 00 in the interior of its arcs, we set $T^{(n+1)}=f_{c}(T^{(n)})$. If instead 00 lies in the interior of one of the arcs of $f_{c}(T^{(n)})$, we replace this arc by $[f_{c}^{n}(x),0]$ and we let $T^{(n+1)}$ be the result of this replacement. By construction, at any step $n$, $T^{(n)}$ is a tripod with 3 arcs and $f_{c}^{n}(x)$ as center. Since each $V_{i}$ is eventually mapped to the attractive cycle $U_{0},U_{1},U_{2}$, for $n$ large enough the ends of $T^{(n)}$ eventually all fall in some $U_{i}$. Since the $U_{0}$ is between $U_{1}$ and $U_{2}$, we have a contradiction with the fact that $T^{(n)}$ is a tripod. This proves point 4 of B. ∎
@@ Appendix C Rabbits and airplane as limit spaces of replacement systems
In this section we show that the basilica, rabbits and airplane limit spaces built with the replacement systems introduced in [BF19] satisfy our A and B thus are homeomorphic to the basilica, rabbits and airplane Julia sets. In order for this section to be as concise as possible, the introduction to replacement systems and limit spaces is reduced to its essential parts and the proofs do not delve into details. We refer to the paper which introduced this topic [BF19] and the second author’s PhD thesis [Tar24a] for a much more thorough introduction to replacement systems, limit spaces and (although we will not use them) rearrangement groups.
@@ C.1.2 The symbol space
Although we do not discuss this here, it is worth recalling that replacement systems allow to easily define countable groups of homeomorphisms of limit spaces, the so-called rearrangement groups. Such groups include Thompson groups $F$, $T$ and $V$ together with new groups acting on fractals such as the basilica [BF15], the airplane [Tar24b] and Ważewski dendrites [Tar23].
@@ C.1.3 Topology of the limit space
    If the base and replacement graphs are connected, then $X$ is connected [BF19, Remark 1.27] and locally connected [Tar24a, Corollary 2.33].
@@ C.2 Arcs and circles in limit spaces
In any directed graph, an undirected walk is a sequence $(e_{1},\dots,e_{k})$ of pairwise adjacent edges. Each undirected walk identifies a sequence of vertices $(v_{0},v_{1},\dots,v_{k})$ that the walk travels through. An undirected walk is an undirected path if there is no $i\neq j$ such that $v_{i}=v_{j}$ and it is an undirected cycle if $v_{i}=v_{j}$ only for $i=0$ and $j=k$.
Remark C.8.
@@
    For each $e_{i}$, choose an unoriented path $(e_{i1},\dots,e_{ik_{i}})$ in its replacement graph joining the initial and terminal vertices.
@@
    Iterate the previous point for each $e_{ij}$.
@@
Instead, in the airplane replacement system (Figure 11) there are infinitely many arcs joining for example the gluing vertices corresponding to the initial and the terminal vertices of a blue edge, since each blue edge requires the choice of one of two red edges and there are infinitely many blue edges at which this choice needs to be made. However, if two distinct vertices belong to a red cycle of some graph expansion, then by Remark C.8 there are precisely two arcs joining them and there is a unique circle that corresponds to that cycle.
@@
The previous remark does not hold unless the procedure described in Remark C.8 stabilizes after finitely many steps. For example, this does not work in the replacement system depicted in Figure 13, whose limit space is homeomorphic to the Julia set for the complex rational map $z\mapsto z^{-2}-1$.

### v2

@@ B.1 The rabbit Julia sets
Let $U_{1}$ be the Fatou component of $c_{1}$ and $U_{0}$ that of the critical point 0. For a component $U$ of $\mathring{K_{c}}$, its image $f(U)$ is also some component of $\mathring{K_{c}}$ (this follows from the fact that $f_{c}$ is an open map) and, if $U$ does not contain 0 (i.e., $U\neq U_{0}$), then $f_{c}$ restricts to a homeomorphism from $U$ to $f_{c}(U)$.
@@
The projection to any component is a cut point. Thus, if $x\in\partial U$ and $y\in\partial V$, where $U$ and $V$ are distinct components of $\mathring{K_{c_{n}}}$, then the projection of $x$ to $V$ is a cut point that separates $x$ from $y$ [DH84, Exposé I, Proposition 4]. This proves point 2 of A
For $x\in K_{c_{n}}$, we denote by $\ell(x)$ the number of components $V$ of $\mathring{K_{c_{n}}}$ such that $V\cap[x,\alpha_{1}]\neq\emptyset$, where $x,\alpha_{1}]$ is the legal arc between $x$ and $\alpha_{1}$ (legal arcs were introduced in [DH84, Exposé II, §6] as arcs réglementaires). By construction of legal arcs, this number is constant on each component $U$ of $\mathring{K_{c}}$, so we denote by $\ell(U)$ this common number (which, a priori, could be infinite). We claim that, for any component $U$, $|\ell(f_{c}(U))-\ell(U)|\leq 1$. If the legal arc between $x\in U$ and $\alpha_{1}$ does not contain 0 then there is equality, since $f_{c}$ maps components to components and, by [DH84, Exposé IV, Lemme 1], it maps legal arcs to legal arcs injectively if the legal arc does not contain the critical point in its interior. In case $x,\alpha_{1}]$ contains 0, we have that $x,\alpha_{1}]=[x,0]\cup[0,\alpha_{1}]$. If $f_{c}(x),f_{c}(0)]=[f_{c}(x),c]$ contains $\alpha_{1}$, then $\ell(f_{c}(U))=\ell(U)-1$ because $f_{c}(x),c]=[f_{c}(x),\alpha_{1}]\cup[\alpha_{1},c]$ and $c$ belongs to the component $U_{1}$ which contains $\alpha_{1}$ in its boundary. If instead $f_{c}(x),c]$ does not contain $\alpha_{1}$, then $\ell(f_{c}(U))=\ell(U)$.
We know that $U_{0}$ contains a point of the attractive cycle, so for any component $U$, there is $n$ such that $f_{c}^{n}(U)=U_{0}$. In particular, $\ell(U)\leq n+1$. So, if $x$ lies on some circle of $J_{c}$, i.e., $x\in\partial U$ for some component $U$ of $\mathring{K_{c}}$, then $\ell(x)=\ell(U)$ or $\ell(x)=\ell(U)-1$, depending on whether $x$ is the projection of $\alpha_{1}$ on $U$ or not. If $x$ and $y$ belong to circles, then the collection of circles that separate $x$ from $y$ is included in the collection of circles that separate $x$ from $\alpha_{1}$ and $y$ from $\alpha_{1}$. This is easily seen using the fact that the legal hull of $x$, $y$ and $\alpha_{1}$ (introduced in [DH84, Exposé II, §7] as enveloppe réglementaire) is a topological tree. This collection is finite, since $\ell(x),\ell(y)<\infty$. This proves point 3 of A.
@@ B.2 The airplane Julia set
Let $\alpha_{1}$ be the point in $J_{c}$ with the external rays $3/7$ and $4/7$. This is the projection of 0 on the component $U_{1}$ of $\mathring{K_{c}}$ that contains $c$. The point $\alpha_{1}$ belongs to the circle $\partial U_{1}$ and is a cut point, since it separates 0 and $c$. The set of its preimages is dense, which proves point 1 of B.
@@
The point $\alpha_{1}$ is mapped to the point with external arguments $\pm 1/7$ which is mapped to the point with external arguments $\pm 2/7$. This last point is mapped to $\alpha_{1}$, which is thus 3-periodic. Let $U_{0}$ be the component that contains 0 and let $U_{2}$ be the image of $U_{1}$. These components form a 3-cycle and $U_{0}$ lies between $U_{1}$ and $U_{2}$. For two components $U,V$ of $\mathring{K_{c}}$, we denote by $\delta(U,V)$ the number of components that the legal arc from any point in $U$ to any point in $V$ crosses minus 1 (this number can be infinite). Since $U_{0}$ is between $U_{1}$ and $U_{2}$, we have $\delta(U_{1},U_{2})=\delta(U_{1},U_{0})+\delta(U_{0},U_{2})$. Because of [DH84, Exposé IV, Lemme 1], we have $\delta(U_{2},U_{1})=\delta(f_{c}(U_{1}),f_{c}(U_{0}))=\delta(U_{1},U_{0})$ and $\delta(U_{0},U_{1})=\delta(f_{c}(U_{2}),f_{c}(U_{0}))=\delta(U_{2},U_{0})$. Since all these numbers are at least $1$, they must all be infinite.
@@
Assume that two circles $C_{1},C_{2}$ have a common point. Let $V_{1},V_{2}$ be the two components of $\mathring{K_{c}}$ such that $C_{i}=\partial V_{i}$. There is a large enough $k\in\mathbf{N}$ such that $f_{c}^{k}$ maps $V_{i}$ to $U_{j_{i}}$ (for $i=1,2$ and $j_{i}\in\{0,1,2\}$) . By [DH84, Exposé IV, Lemme 1], $f_{c}^{k}$ maps the legal arc between $V_{1}$ and $V_{2}$ to the legal arc between $U_{j_{1}}$ and $U_{j_{2}}$, so it maps the point shared by $V_{1}$ and $V_{2}$ to a point shared by two of the $U_{i}$’s. We have a contradiction, since the $U_{i}$’s have at least one circle that separates them (in fact, infinitely many). This proves point 3 of B.
@@
Assume that two circles $C_{1},C_{2}$ are not separated by a third one. Let $V_{i}$’s be components as above. The legal arc $A$ joining the centers does not contain 0 and the same is true for all $f_{c}^{n}(A)$. Let $n$ be a large enough number so that $f^{n}_{c}(V_{i})$’s are some $U_{j}$’s. The images of $V_{1}$ and $V_{2}$ are distinct and thus separated by infinitely many circles. This implies that $A$ meets infinitely many circles. A contradiction, which proves point 2 of B.
@@
It remains to prove that every cut point has order $2$. Since every cut point has order at least $2$, it suffices to prove there is no cut point of order at least $3$. By contradiction, let $x$ be such a point. Since $K_{c}$ is obtained by filling circles from $J_{c}$, the complement $K_{c}\setminus\{x\}$ has at least $3$ connected components. Choose components $V_{1},V_{2},V_{3}$ of $\mathring{K_{c}}$ lying in distinct components of $K_{c}\setminus\{x\}$. Let $T$ the legal hull of the centers of the $V_{i}$’s. Up to choosing a smaller tree, we may assume that $T$ does not contain 0 in the interior of any of its arcs. It is a tripod with $3$ arcs and $x$ as center. We define a sequence of legal trees in the following way. Let $T^{(0)}=T$ and assume that $T^{(n)}$ has been constructed and does not contain 0 in the interior of any of its arcs. Then, by [DH84, Exposé IV, Lemme 1], $f_{c}$ maps $T^{(n)}$ injectively to $f_{c}(T^{(n)})$. If the latter does not contain 0 in the interior of its arcs, we set $T^{(n+1)}=f_{c}(T^{(n)})$. If instead 0 lies in the interior of one of the arcs of $f_{c}(T^{(n)})$, we replace this arc by $f_{c}^{n}(x),0]$ and we let $T^{(n+1)}$ be the result of this replacement. By construction, at any step $n$, $T^{(n)}$ is a tripod with 3 arcs and $f_{c}^{n}(x)$ as center. Since each $V_{i}$ is eventually mapped to the attractive cycle $U_{0},U_{1},U_{2}$, for $n$ large enough the ends of $T^{(n)}$ eventually all fall in some $U_{i}$. Since the $U_{0}$ is between $U_{1}$ and $U_{2}$, we have a contradiction with the fact that $T^{(n)}$ is a tripod. This proves point 4 of B. ∎
B.3 Extending homeomorphisms to the complex plane
Since Julia sets lie by definition in the complex plane, it is natural to determine which homeomorphisms of a Julia extend to a global homeomorphism of the complex plane. Let us show that such homeomorphisms correspond to the elements of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ from Section 3.
We denote by $\mathrm{Homeo}^{+}_{J}(\mathbf{C})$ the group of orientation preserving homeomorphisms of $\mathbf{C}$ stabilizing $J$ setwise, i.e. $g(J)=J$.
Theorem B.5.
Let $J$ be the airplane Julia set or a rabbit Julia set. The subgroup of $\mathrm{Homeo}(J)$ of homeomorphisms that are the restrictions of elements of $\mathrm{Homeo}^{+}_{J}(\mathbf{C})$ coincide with $\mathrm{Homeo}^{+}(J)$.
Proof.
Let $g\in\mathrm{Homeo}(J)$ that is the restriction to $J$ of some $\overline{g}\in\mathrm{Homeo}^{+}_{J}(\mathbf{C})$. Since the Riemann sphere $\hat{\mathbf{C}}$ is the Alexandroff compactification of $\mathbf{C}$, $\overline{g}$ extends uniquely as a homeomorphisme of $\hat{\mathbf{C}}$ fixing $\infty$. Take three points $x,y,z$ on some circle $C$ of $J$ and arcs $I_{x},I_{y},I_{z}$ such that $\infty$ is one end and $x$, respectively $y$ or $z$, is the other end, and such that $I_{x}\cap J=\{x\}$, respectively $\{y\}$ or $\{z\}$. We may moreover assume that these arcs intersect only at $\infty$. For example, external rays do the job.
Since $\overline{g}$ is a homeomorphism of $\hat{\mathbf{C}}$, the arcs $\overline{g}(I_{x}),\overline{g}(I_{y})$ and $\overline{g}(I_{z})$ satisfy the same conditions and thus $O^{C}_{\varphi}(x,y,z)=O^{g(C)}_{\varphi}(g(x),g(y),g(z))$. So $g\in\mathrm{Homeo}^{+}(J)$.
Conversely, consider $g\in\mathrm{Homeo}^{+}(J)$. Then there is a $\gamma\in\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ preserving the lamination $\mathcal{L}J$ such that $g\circ\varphi=\varphi\circ\gamma$. Let $\mathbf{D}$ be the closed unit disk in $\mathbf{C}$ and $\Phi\colon\hat{\mathbf{C}}\setminus\mathbf{D}\to\hat{\mathbf{C}}\setminus FJ$ be the Böttcher coordinates of the complement of the filled in Julia set $FJ$. In particular, for any $z$ in the unit circle $\mathbf{S}^{1}$ of $\mathbf{C}$, $\varphi(z)=\lim_{r\to 1}\Phi(rz)$ and this defines a continuous extension of $\Phi$ to the complement of the open unit disk. We define $\overline{g}$ on $\hat{\mathbf{C}}\setminus FJ$ ($r>1$ and $z\in\mathbf{S}^{1}$) via the formula
$\overline{g}(\Phi(rz))=\Phi(r\gamma(z)).$
This defines a homeomorphism of $(\hat{\mathbf{C}}\setminus FJ)\cup J$. Any component $U$ of the interior of $FJ$ is homeomorphic to an open disk with boundary $\partial U$ which is a circle of $J$. For such $U$ we choose a homeomorphism between $U$ and the component $V$ of the interior of $FJ$ bounded by $g(\partial U)$ which extends to a homeomorphism between $\overline{U}$ and $\overline{V}$ that coincides with $g$ on $\partial U$. Patching together all these homeomorphisms of interior components we get a global homeomorphism $\overline{g}$ of $\hat{\mathbf{C}}$ whose restriction to $J$ is $g$. ∎
Remark B.6.
The construction of the extension done in the above proof is not unique at all, but one can make it canonical. For each component $U$ of the filled-in Julia center, one can choose a center $c_{U}$ and this defines internal rays and arguments. Doing a similar construction to the one we do for external rays, one can define uniquely $\overline{g}$ on each component $U$ such that it maps internal rays of $U$ to internal rays of $V$. This way, one can see that $\mathrm{Homeo}^{+}_{J}(\mathbf{C})$ splits as permutational wreath product
$\left(\mathrm{Homeo}_{0}(\mathbf{C}\setminus FJ)\times\prod_{U\in\pi_{0}(\overset{\circ}{FJ})}\mathrm{Homeo}_{0}(U)\right)\rtimes\mathrm{Homeo}^{+}(J),$
where $\mathrm{Homeo}_{0}(U)$ is the subgroup of $\mathrm{Homeo}(U)$ that admits a continuous extension to $\overline{U}$ that is the identity on $\partial U$.
@@ Appendix C Rabbits and airplane as limit spaces of replacement systems
In this section we show that the basilica, rabbits and airplane limit spaces built with the replacement systems introduced in [BF19] satisfy our A and B thus are homeomorphic to the basilica, rabbits and airplane Julia sets. In order for this section to be as concise as possible, the introduction to replacement systems and limit spaces is reduced to its essential parts and the proofs do not delve into details. We refer to the paper which introduced this topic [BF19] and the second author’s PhD thesis [Tar24] for a much more thorough introduction to replacement systems, limit spaces and (although we will not use them) rearrangement groups.
@@ C.1.2 The symbol space
Although we do not discuss this here, it is worth recalling that replacement systems allow to easily define countable groups of homeomorphisms of limit spaces, the so-called rearrangement groups. Such groups include Thompson groups $F$, $T$ and $V$ together with new groups acting on fractals such as the basilica [BF15], the airplane [Tar24a] and Ważewski dendrites [Tar25].
@@ C.1.3 Topology of the limit space
    If the base and replacement graphs are connected, then $X$ is connected [BF19, Remark 1.27] and locally connected [Tar24, Corollary 2.33].
@@ C.2 Arcs and circles in limit spaces
In a directed graph, an undirected walk is a sequence $(e_{1},\dots,e_{k})$ of edges such that $e_{i}$ is adjacent to $e_{i+1}$ for all $i\in\{1,\dots,k-1\}$. Each undirected walk identifies a sequence of vertices $(v_{0},v_{1},\dots,v_{k})$ that the walk travels through. An undirected walk is an undirected path if there is no $i\neq j$ such that $v_{i}=v_{j}$ and it is an undirected cycle if $v_{i}=v_{j}$ only for $i=0$ and $j=k$.
Proposition C.8.
@@
    For each edge $e_{i}$, choose an unoriented path $(e_{i1},\dots,e_{ik_{i}})$ in its replacement graph joining the initial and terminal vertices.
@@
    Iterate the previous point for each edge $e_{ij}$.
@@
Proof.
First note that, for each graph expansion featuring $p$ and $q$ as vertices, an arc joining $p$ and $q$ in the limit space is included in the union of cells corresponding to a unique unoriented path between $p$ and $q$ (this is because non-empty intersections of cells consist of one or two gluing vertices, so intervals can only travel between cells corresponding to adjacent edges by going through such gluing vertices). By Proposition 1.24 of [BF19] a cell associated to an edge $e_{1}$ is included in one associated to an edge $e_{2}$ if and only if $e_{2}$ is a prefix of $e_{1}$ (as finite words). Thus, given a graph $\Gamma_{1}$ that features $p$ and $q$ as vertices and a graph $\Gamma_{2}$ that is an expansion of $\Gamma_{1}$, if an arc joining $p$ and $q$ is included in the union of cells of unoriented paths $P_{1}$ in $\Gamma_{1}$ and $P_{2}$ in $\Gamma_{2}$, then $P_{2}$ must be obtained as an expansion of edges of $P_{1}$. Thus, every arc is as in the statement.
For the converse, the idea is that the gluing relation in a sequence of “nested” paths as in the statement behaves like a generalization of $n$-ary expansions of numbers in $[0,1]$ with no fixed arity (see [BF19], where Proposition 2.6 explains how to obtain the interval as a limit space with dyadic numbers as gluing vertices and Example 2.14 features examples whose limit spaces are intervals and the gluing vertices have no fixed arity). Consider a sequence $P_{1},P_{2},\dots$ of paths as in the statement. Let $Y$ be the set of classes of infinite words whose every long enough prefix, as an edge, lies in some $P_{i}$. By [BF19, Proposition 1.22], two elements of the symbol space are equivalent under the gluing relation if and only if they represent the same gluing vertex. Thus, for any $x=x_{1}x_{2}\dots$ and $y=y_{1}y_{2}\dots$ in the symbol space whose all long enough prefixes lie in some $P_{i}$, one has that $x\sim y$ if and only if all the long enough prefixes $x_{1}\dots x_{k}$ and $y_{1}\dots y_{k}$ are adjacent edges. Since these prefixes all lie in the paths $P_{i}$’s, the gluing relation on $Y$ is non-trivial precisely on any vertex on the $P_{i}$’s, where it glues the two sequences of edges approaching from the two sides (compare with the binary gluing relation $w0111\ldots\sim w1000\ldots$ described in [BF19, Proposition 2.6]). Hence $Y$ is this homeomorphic to $[0,1]$ if $p\neq q$ and to $\mathbf{S}^{1}$ if $p=q$. ∎
@@
Instead, in the airplane replacement system (Figure 11) there are infinitely many arcs joining for example the gluing vertices corresponding to the initial and the terminal vertices of a blue edge, since each blue edge requires the choice of one of two red edges and there are infinitely many blue edges at which this choice needs to be made. However, if two distinct vertices belong to a red cycle of some graph expansion, then by Proposition C.8 there are precisely two arcs joining them and there is a unique circle that corresponds to that cycle.
@@
The previous remark does not hold unless the procedure described in Proposition C.8 stabilizes after finitely many steps. For example, this does not work in the replacement system depicted in Figure 13, whose limit space is homeomorphic to the Julia set for the complex rational map $z\mapsto z^{-2}-1$.